If real GDP grows at an annual rate of 1.49% then we can expect real GDP to double in approximately how many years? Enter a number rounded to two decimal places. Please show work
Answer #1

Answer

By rule of the 70

years to double =70/g

=70/1.49

=46.9798658

=46.98 years

it will tak 46.98 years approximately

==========

By compounding formula

years =ln(2)/ln(1+i)

years=ln(2)/ln(1+0.0149)

=46.8656643

=46.87 years

It will take 46.87 years